# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

### v3

@@ Abstract
Magnetar is a neutron star with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G) which usually manifests as soft gamma-ray repeater (SGR) or anomalous X-ray pulsar (AXP). SGR J1935+2154 is not only one of the most active magnetar detected so far, but also the unique confirmed source of fast radio burst (FRB). Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM) are dedicated to monitor gamma-ray transients all over the sky, including SGR bursts. Here we report the GECAM observation of the burst activity of SGR J1935+2154 from January 2021 to December 2022, which results in a unique and valuable data set for this important magnetar. With a targeted search of GECAM data, 164 bursts from SGR J1935+2154 are detected by GECAM-B while 97 bursts by GECAM-C, including the X-ray burst associated with a fast radio burst (FRB 20221014). We find that both the burst duration and the waiting time between two successive bursts follow lognormal distributions. The period of burst activity is $134\pm 20$ days, thus the burst activity could be generally divided into 4 active episodes over these two years. Interestingly, the hardness ratio of X-ray bursts tends to be softer and more concentrated over these two years, especially during the active episode with FRBs detected.
@@ 1 Introduction
The soft gamma-ray repeaters (SGRs) and anomalous X-ray pulsars (AXPs) are perceived to originate in the highly magnetized neutron star, namely magnetar (Duncan & Thompson, 1992; van Kerkwijk et al., 1995; Kouveliotou et al., 1998; Banas et al., 1997; Kaspi & Beloborodov, 2017). Magnetars are characterized by slow rotation period ($P\sim 2-12\ \mathrm{s}$), rapidly spin down ($\dot{P}\sim 10^{-13}-10^{-11}\ \mathrm{s\cdot s^{-1}}$) and relatively young age (typically about 1000 yr). So far, 30 magnetars had been detected and 24 of them had been confirmed (Olausen & Kaspi, 2014). Most magnetars can increase persistent radiation and emit bursts/flares simultaneously in the X-/gamma-ray band during an outburst.
@@
Based on their luminosity and duration, the SGR bursts can be divided into three classes (Woods & Thompson, 2006): Short-duration burst, which consists of single or multiple pulses, is the most typical magnetar burst with a burst duration $\sim$0.01 $\mathrm{s}$ and the luminosity around $10^{38}\sim 10^{40}$ $\mathrm{erg\cdot s^{-1}}$; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ($>1$ $\mathrm{s}$) and peak luminosity around $10^{41}\sim 10^{43}$ $\mathrm{erg\cdot s^{-1}}$; Giant flare, the rarest and the most powerful energetic burst, is characterized by a significantly higher luminosity than a typical magnetar burst and a special pulse profile with a hard initial spike and rapidly decaying tail.
@@
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos et al., 2014), while it is worth to note that Xie et al. (2022) found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan et al., 2009) before the this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, $P\sim 3.24\ \mathrm{s}$ and $\dot{P}\sim 1.43\times 10^{-11}\ \mathrm{s\cdot s^{-1}}$, respectively. This indicates a dipole-magnetic field of $B\sim 2.2\times 10^{14}\ \mathrm{G}$ (Israel et al., 2016). It has experienced multiple active episodes from 2014 to 2022 (Younes et al., 2017; Lin et al., 2020b, a; Cai et al., 2022a, b; Xie et al., 2022; Rehan & Ibrahim, 2023, 2024). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which a burst forest was observed (Kaneko et al., 2021). Among these bursts there is an X-ray burst (XRB) (Li et al., 2021; Mereghetti et al., 2020; Ridnaia et al., 2021; Tavani et al., 2020) associated with a fast radio burst, FRB 200428 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020a). After the FRB 20200428 event, GECAM (Li et al., 2021; Xiao et al., 2022a; Zhang et al., 2023; Wang et al., 2024) detected an X-ray burst from SGR J1935+2154 which associated with a radio burst detected by CHIME on 2022 October 14 (Wang et al., 2022a). Later, the Insight–Hard X-ray Modulation Telescope (HXMT) again detected an X-ray burst associated with a radio burst found by the Yunnan 40 m telescope on 2022 October 21 (Li et al., 2022).
@@
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observation data from January 2021 to December 2022. As Paper I of this series work for GECAM observation of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (In prep.), we will give the detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section 2, we present the observation, search process, and report the burst catalog. In Section 3, we present the catalog analysis method and results. Finally, the summary and discussions are given in Section 4.
@@ 2 Observation and Burst Search
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, Xiao et al., 2022b), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang et al., 2023) and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang et al., 2024), dedicated to monitor all-sky gamma-ray transients. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 Charged particle detectors (CPDs) while GECAM-C has 12 GRDs and 2 CPDs. All GRDs of GECAM-B and most of GRDs of GECAM-C operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time.
@@ 2.1 Search and Identification
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan & Ibrahim, 2023, 2024; Wood & Fermi-GBM Team, 2022). A search pipeline developed by Cai et al. (2021) for gamma-ray bursts (GRBs) is used to do a blind search on GECAM or HXMT daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B, as reported in Xie et al. (2022) and 75 X-ray bursts were found by HXMT (Cai et al., 2022a).
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold, we conduct a targeted sub-threshold search on GECAM daily event-by-event (EVT) data [^1] https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai et al., submitted. Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin et al., 2020b, a; Rehan & Ibrahim, 2023, 2024), we utilize the following spectral templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameters setting is listed in Table 1 which is similar with the Table 1 in Xie et al. (2022). The details of each template are presented below.
@@
Thousands of trigger (burst candidates) have been found through the targeted search process. We execute a refined analysis for each burst candidate by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools [^2] https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from candidates should adhere to the following criteria (same as Xie et al., 2022):
@@
    Using the GECAM position history file (posatt) to retrieve the position of satellite and the pointing of detector. The candidates are excluded when SGR J1935+2154 is eclipsed by Earth;
@@
    The candidates with the angle of the three most significant detectors relative to the source position (R.A. = 293.73°, decl. = 21.90°, Israel et al., 2016) should be less than $60^{\circ}$ (with at least two detectors);
@@
    The solar flares, the particle events measured by CPD observation, and the Terrestrial gamma-ray flash (TGF, Briggs et al., 2013) and Terrestrial electron beam (TEB, Xiong et al., 2012) featured by being harder ($\sim$ MeV) and shorter ($\sim$ 1 ms) than magnetar bursts are excluded;
@@
    Identified by other instruments (e.g. Fermi/GBM).
@@
GECAM-B SoftBand 27 16.46%
CPL 39 23.78%
OTTB 57 34.75%
Blackbody 31 18.90%
Powerlaw 10 6.09%
@@
All bursts SoftBand 34 12.84%
CPL 74 17.15%
OTTB 81 23.04%
Blackbody 50 15.12%
Powerlaw 24 7.85%
@@ 2.2 Search Results
A total of 164 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a fast radio burst FRB 20221014 (Wang et al., 2022b). It is worth to note that about one third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM (blocked by the Earth or turned off during the SAA passage).
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to targeted search. The CPL, OTTB, and BB account for a proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai et al., 2021). As shown in Fig. 3, Using a Powerlaw template to search magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing much difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel (¿40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to found using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
@@
Burst ID Trigger Time Template R.A. Decl. Err Sig.aaSignificance obtained by search pipeline. $T_{\mathrm{bb}}$bbBurst duration derived by a Bayesian block method. $T_{\mathrm{st}}$ccBurst start time relative to the trigger time. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBMddCalculated with the Fermi/GBM position history file (poshist) using the GBM data tool (Goldstein et al., 2022). GECAM-C
@@
130 2022-10-12T22:34:56.900 SoftBand 299.32 5.40 9.37 7.50 29.84 -43.19 No No
131 2022-10-13T00:00:53.450 SoftBand 291.38 20.32 1.10 8.41 348.81 -92.94 Yes No
132 2022-10-14T06:47:24.100 SoftBand 299.64 22.19 5.12 13.70 242.99 -150.79 No Yes
133 2022-10-14T07:12:28.800 SoftBand 300.00 19.45 3.44 89.16 214.46 -74.52 Yes Yes
@@
Burst ID Trigger Time Template R.A. Decl. Err Sig.aaSignificance obtained by search pipeline. $T_{\mathrm{bb}}$bbBurst duration derived by a Bayesian block method. $T_{\mathrm{st}}$ccBurst start time relative to the trigger time. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBMddCalculated with the Fermi/GBM position history file (poshist) using the GBM data tool (Goldstein et al., 2022). GECAM-B

### v5

@@ Abstract
Magnetar is a neutron star with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G). The magnetar SGR J1935+2154 is not only one of the most active magnetars detected so far, but also the unique confirmed source of fast radio bursts (FRBs). Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM) is dedicated to monitor gamma-ray transients all over the sky, including magnetar short bursts. Here we report the GECAM observations of the burst activity of SGR J1935+2154 from January 2021 to December 2022, which results in a unique and valuable data set for this important magnetar. With a targeted search of GECAM data, 159 bursts from SGR J1935+2154 are detected by GECAM-B while 97 bursts by GECAM-C, including the X-ray burst associated with a bright radio burst. We find that both the burst duration and the waiting time between two successive bursts follow lognormal distributions. The period of burst activity is $134\pm 20$ days, thus the burst activity could be generally divided into four active episodes over these two years. Interestingly, the hardness ratio of X-ray bursts tends to be softer during these two years, especially during the active episode with radio bursts detected.
@@ 1 Introduction
Magnetars are highly magnetized isolated neutron stars (Duncan & Thompson, 1992; van Kerkwijk et al., 1995; Kouveliotou et al., 1998; Banas et al., 1997; Kaspi & Beloborodov, 2017), and are characterized by their slow rotation period (P212similar-toabsent212$\sim$ 2-12∼ 2 - 12 s), rapidly spinning down ($\dot{P}\sim 10^{-13}-10^{-11}$ $\mathrm{s\cdot s^{-1}}$), and relatively young age (typically about 1000 yr). We currently know about 30 magnetars (Olausen & Kaspi, 2014) [^1] http://www.physics.mcgill.ca/~pulsar/magnetar/main.html. Magnetars are manifested themselves by persistent emission and emit bursts/flares simultaneously in the X-/gamma-ray band (Esposito et al., 2021).
@@
Based on their luminosity and duration, magnetar bursts can be divided into three classes (Woods & Thompson, 2006): Short-duration burst, which consists of single or multiple peaks, is the most typical magnetar burst with a burst duration $\sim$0.01 $\mathrm{s}$ and the luminosity around $10^{38}\sim 10^{40}$ $\mathrm{erg\cdot s^{-1}}$; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ($>1$ $\mathrm{s}$) and peak luminosity around $10^{41}\sim 10^{43}$ $\mathrm{erg\cdot s^{-1}}$; Giant flare, the rarest and the most powerful energetic burst is characterized by a significantly higher luminosity than a typical magnetar burst and a light curve with a hard initial spike and rapidly decaying tail modulated at the spin period of the magnetar.
@@
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by the Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos et al., 2014), while it is worth noting that Xie et al. (2022) found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan et al., 2009) before this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, $P\sim 3.24\ \mathrm{s}$ and $\dot{P}\sim 1.43\times 10^{-11}\ \mathrm{s\cdot s^{-1}}$, respectively. This indicates the estimated magnetic field at the pole is $B\sim 2.2\times 10^{14}\ \mathrm{G}$ (Israel et al., 2016). It has experienced multiple outbursts from 2014 to 2022 (Younes et al., 2017; Lin et al., 2020b, a; Cai et al., 2022a, b; Xie et al., 2022; Rehan & Ibrahim, 2023, 2024). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which burst forests were observed (Kaneko et al., 2021; Younes et al., 2020). Moreover, the source entered a new outburst (Borghese et al., 2022). Among these bursts there is an X-ray burst (Li et al., 2021; Mereghetti et al., 2020; Ridnaia et al., 2021; Tavani et al., 2020) associated with a fast radio burst, FRB 200428 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020a). Additional coincident radio and X-ray bursts were observed during the active episode of 2022 (Li et al., 2021; Xiao et al., 2022a; Zhang et al., 2023; Wang et al., 2024, 2022a; Li et al., 2022).
@@
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observations data from January 2021 to December 2022. As in Paper I of this series work for GECAM observations of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (in prep.), we will give a detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section 2, we present the observations, the search process, and report the burst catalog. In Section 3, we present the catalog analysis method and the results. Finally, a summary and discussions are given in Section 4.
@@ 2 Observation and Burst Search
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, Xiao et al., 2022b), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang et al., 2023), and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang et al., 2024). These four satellites are all-sky monitors with a large field of view that could monitor various gamma-ray transients except for the area blocked by Earth. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 charged particle detectors (CPDs), while GECAM-C has 12 GRDs and 2 CPDs. All GECAM-B’s GRDs and most GECAM-C’s GRDs operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The event-by-event (EVT) data from GECAM’s GRD detectors is used for further analysis. The GECAM achieves a high time resolution of 0.1 $\mathrm{\mu s}$, with a dead time of 4 $\mathrm{\mu s}$ for normal events and exceeding 69 $\mathrm{\mu s}$ for overflow events (Liu et al., 2021). We note that in this work only GECAM-B/C data are used because of power supply issues that GECAM-A has not been able to observe the source yet (Li et al., 2022).
@@ 2.1 Search and Identification
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan & Ibrahim, 2023, 2024; Wood & Fermi-GBM Team, 2022). A search pipeline developed by Cai et al. (2025) for gamma-ray bursts (GRBs) is used to do a blind search on GECAM daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B (2021-01 to 2022-01, Xie et al., 2022) [^2] The 94 bursts shown in Xie et al. (2022) are the preliminary blind searched results that are reported in the Gamma-ray Coordinates Network (GCN) by the burst advocate (BA). We identify 82 bursts from SGR J1935+2154 using the targeted search in this paper..
The search pipeline is implemented using a ground sensitive coherent search method that features a log-likelihood ratio (hereafter LR) comparing expected model counts to observed counts in each channel of each detector (Blackburn et al., 2015; Cai et al., 2025),
$\mathcal{L}=\sum_{i=1}^{j}[\ln\frac{\sigma_{n_{\mathrm{i}}}}{\sigma_{d_{\mathrm{i}}}}+\frac{\widetilde{d_{\mathrm{i}}^{2}}}{2\sigma_{n_{\mathrm{i}}}^{2}}-\frac{(\widetilde{d_{\mathrm{i}}}-r_{\mathrm{i}}s)^{2}}{2\sigma_{d_{\mathrm{i}}}^{2}}],$ (1)
where $i$ is the number of data sets in each channel and detector, $j$ is the total number of detectors and channels, $\widetilde{d_{\mathrm{i}}^{2}}$ is the background-subtracted counts and $\sigma_{d_{\mathrm{i}}}$ is the error of the expected counts (background+source), $n_{\mathrm{i}}$ is the estimated background and $\sigma_{n_{\mathrm{i}}}$ is the error of the background counts, $r_{\mathrm{i}}$ and $s$ represent the instrument response and the expected source amplitude (could estimated by a spectral model template), respectively. Therefore, the LR could be calculated with the response factors over the all-sky for a blind search or a certain area for a targeted search. Then, the continuous time-bins with the higher LR (higher than the LR threshold) could be considered as one signal (trigger/burst). The maximum LR gives the most possible spatial location of the source signal. The pipeline will mark the first time-bin, of which LR is higher than the LR threshold, as a trigger time. In this paper, the LR threshold is set to 20 and the search timescale (time-bin) is set to 10 ms.
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold (typically $5\sigma$ as implemented in flight software), we conduct a targeted sub-threshold search for the area around the position of SGR J1935+2154 (R.A. and Decl. ±$\pm 20^{\circ}$) on GECAM daily EVT data [^3] https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai et al. (2025). Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral model templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin et al., 2020b, a; Rehan & Ibrahim, 2023, 2024), we utilize the following spectral model templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameter setting is listed in Table 1 which is similar to Table 1 in Xie et al. (2022). The details of each template are presented below.
@@
Thousands of triggers (candidate bursts) have been found through the targeted search process. We perform a refined analysis for each candidate burst by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools [^4] https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from the candidates should adhere to the following criteria (same as Xie et al., 2022):
@@
    Using the GECAM position history file (posatt) to retrieve the position of satellite and the pointing of detector. The candidates are excluded when SGR J1935+2154 is occulted by Earth;
@@
    The burst location is within the error circle (at 3 $\sigma$) of the source position (SGR J1935+2154, R.A. = 293.73°, Decl. = 21.90°, Israel et al., 2016);
@@
    The solar flares, the particle events measured by CPD observations, and the Terrestrial gamma-ray flash (TGF, Briggs et al., 2013) and Terrestrial electron beam (TEB, Xiong et al., 2012) featured by being harder ($\sim$ MeV) and shorter ($\sim$ 1 ms) than magnetar bursts are excluded;
@@
    The pipeline will calculate the significance [^5] In this paper, the significance is defined as $S/\sqrt{B}$, where $S$ is the source (background-subtracted) counts and $B$ is the estimated background counts. of each detector. At least two most significant detectors should have an angle relative to the source position of less than $60^{\circ}$ or maybe a false trigger (e.g., particle events);
  5. 5.
    Identified by other instruments (e.g., Fermi/GBM).
@@
GECAM-B SoftBand 27 16.98%
CPL 39 24.52%
OTTB 55 34.59%
Blackbody 29 18.23%
Powerlaw 9 5.66%
@@
All bursts SoftBand 34 24.19%
CPL 74 60.60%
OTTB 79 59.33%
Blackbody 48 37.81%
Powerlaw 21 18.03%
@@ 2.2 Search Results
A total of 159 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a bright radio burst FRB 20221014 (Dong & Chime/Frb Collaboration, 2022; Wang et al., 2022b; Frederiks et al., 2022; Maan et al., 2022; Giri et al., 2023) [^6] We marker this radio burst as FRB 20221014 in this paper since the luminosity is several orders higher than the pulsed emission from SGR J1935+2154 (Zhu et al., 2023) although still a few orders lower than that of general FRBs.. It is worth noting that about one-third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM [^7] Using the Fermi/GBM position history file (poshist) and the GBM data tool (Goldstein et al., 2022) to determine if SGR J1935+2154 is visible to Fermi/GBM or occulted by the Earth at these burst times..
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to the targeted search. The CPL, OTTB, and BB account for a proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai et al., 2021). As shown in Fig. 3, using a Powerlaw template to search for a magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing a much difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel ($>$40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to found using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
@@
Burst ID Trigger Time Template R.A. Decl. Err Sig. $T_{\mathrm{bb}}$aaBurst duration derived by the Bayesian blocks algorithm. $T_{\mathrm{st}}$bbBurst start time relative to the trigger time using the Bayesian blocks algorithm. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBM GECAM-C
@@
127 2022-10-12T22:34:56.900 SoftBand 299.32 5.40 9.37 7.50 29.84 -43.19 No Yes
128 2022-10-13T00:00:53.450 SoftBand 291.38 20.32 1.10 8.41 348.81 -92.94 Yes Yes
129 2022-10-14T06:47:24.100 SoftBand 299.64 22.19 5.12 13.70 242.99 -150.79 No No
130 2022-10-14T07:12:28.800 SoftBand 300.00 19.45 3.44 89.16 214.46 -74.52 Yes No
@@
Burst ID Trigger Time Template R.A. Decl. Err Sig. $T_{\mathrm{bb}}$aaBurst duration derived by the Bayesian blocks algorithm. $T_{\mathrm{st}}$bbBurst start time relative to the trigger time using the Bayesian blocks algorithm. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBM GECAM-B